OpenStudy (anonymous):

the percentage by mass of elements in a particular pure substance will vary from sample to sample, True or false

OpenStudy (anonymous):

lets say you have water and peroxide as example where water is H2O and peroxide is H2O2 now lets simplify it and tell hydrogen atom weighs 10 g and oxygen weighs 20g lets see percentages H2O -> 2 H and 1 O = 2*10g + 20 g = 40 g and percent of H in that compound is 2*10/40 = 0,5 or 50% now for H2O2 -> 2 H and 2 O = 2*10g + 2*20g = 60g and percent of H in that compound is 2*10/60 = 0,33 = 33% so what do you think about your question after knowing this?